OpenStudy (anonymous):

find all solutions in the interval [0,2pi] for the equation 2cos^2 feta-1=0

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

cos^2 (theta) = 1/2 cos (theta) = 1/root2, -1/root2 theta = pi/4, 3pi/4, 5pi/4, 7pi/4

OpenStudy (anonymous):

thanks